40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 5 09:32:12 2012, steve, Update, SUS, PRM oscillation  PRMosc.png
    Reply  Wed Jun 6 09:27:09 2012, steve, Update, SUS, PRM damping restored  PRMwhat.png
       Reply  Thu Jun 7 08:12:20 2012, steve, Update, SUS, PRM is still oscillating PRMoscil.png
          Reply  Thu Jun 7 13:18:42 2012, rana, Update, SUS, PRM is still oscillating 
             Reply  Thu Jun 7 13:25:37 2012, Koji, Update, SUS, PRM is still oscillating 
                Reply  Thu Jun 7 18:46:13 2012, yuta, Update, SUS, PRM oplev centered 
Message ID: 6786     Entry time: Thu Jun 7 13:25:37 2012     In reply to: 6785     Reply to this: 6788
Author: Koji 
Type: Update 
Category: SUS 
Subject: PRM is still oscillating 

Done.

C1:SUS-PRM_OLPIT_GAIN 1.0 -> 0
C1:SUS-PRM_OLYAW_GAIN -0.7 -> 0
 

Quote:

  Set the PRM OL servo gains to zero until someone can take care of this. Turning off the buttons doesn't help anything if people run the alignment scripts.

 

ELOG V3.1.3-